OpenStudy (anonymous):

how can i figure out how to graph -x+3y=6 and 3x-9y=9

OpenStudy (anonymous):

-x + 3y = 6 divde -x to cancel out -x, to make it 3y = -x +6, then take 3y and divide it to make y = -x + 3 I think, I might be wrong but thats how I did it in Algebra I & II